A company had earnings per share of $3.32, a Return on Equity of 12%, and paid quarterly dividends of $0.13 on each share over the last year. You think that the company is likely to use the same payout ratio in the future. Assuming that the company can sustain its current R.O.E. forever, what is the rate at which it can annually grow its dividends forever? Go out three decimals - for example, type 5.1% as .051.
Answer #1

Annual Dividend = Quarterly Dividend * No. of quarters in a year = $0.13 * 4 = $0.52

Payout Ratio = Dividend / EPS = $0.52 / $3.32 = 0.1566, or 15.66%

SGR = ROE * (1 - Payout Ratio)

= 12% * (1 - 0.1566) = 10.12%, or 0.101